A company is interested in examining elasticity of a new product. Economists at that company estimate that short run and long run demands are QD=400-4P and QD= 200-1P respectively. At price of P=75

  1. What is the price elasticity of demand in the short run?

  2. What is the price elasticity of demand in the long run?

  3. Given your answers in 1) and 2), is this product a durable or non-durable good?
Answer #1

When P=75

QD in the short run = 400-4*75 = 100

QD in the long run = 200-1*75 = 125

Elasticity in the short run = -4*75/100 = -3

Elasticity in the long run = -1*75/125 = -0.6

The product is a durable good as the price increase will cause the demand to fall in the long run